Hi all, i need some help with the following. I would like to know how to do this and similiar problems so please show all working with steps and explanations where possible. I need to solve for x for the following:
1/5x + 1/x = 1/15
Thanks, -nick.

Answer by RAY100(1637)

1/(5x) +1/x =1/15
.
LCD = 15x, therefore multiply by 15x
.
3 +15 =x
.
18 =x,,,,,,,answer
.
checking
.
1/(5*18) + 1/18 =1/15
.
.011 + .055 = .066
.
.066 - .066,,,,,,ok